OpenStudy (anonymous):

What is the simplified form of: 7th root of w^28 x^7/y^14

OpenStudy (anonymous):

divide each exponent by 7

OpenStudy (id21):

\[\sqrt[7]{\frac{w^{28}x^7}{y^{14}}}=\frac{w^{\frac{28}{7}}x^{\frac{7}{7}}}{y^{\frac{14}{7}}}=\frac{w^4x}{y^2}\]

OpenStudy (anonymous):

So....... w^4 x/y^2??????

OpenStudy (anonymous):

THANK YOU!!!!!!!!!!!!!

OpenStudy (id21):

:)

OpenStudy (anonymous):

:-)